OpenStudy (tatumlee):

the width of bens rectangular family room is 3 ft less than the length. the perimiter is 70 ft. which equation can be used to find each length of the l of the room? how do I solve this?

OpenStudy (anonymous):

You would start with saying "x" or whatever variable is the length, then saying "x-3" for the width, and knowing that the formula for the perimeter of a rectangle is 2L + 2W, you would say 2x + 2(x-3) = 70 and solve from there. Good luck!

OpenStudy (tatumlee):

so would the answer be 70=2(l-3) ? @Catharsic

OpenStudy (anonymous):

Hmm, that doesn't look quite right. "2x + 2(x-3)" would translate to "4x - 6" after you solved it. That answer leaves out "2x". (or 4l - 6 or whatever you wanna call the variable)

OpenStudy (tatumlee):

I really am just confused, the answer options are: a. 70= l-3, b.70=2l-3, c. 70=2(l-3), d. 70= 2(2l-3)

OpenStudy (anonymous):

Yep, one of those is right. Solve "2l + 2(l-3)" and then find what that is (4l - 6). Look at those options - which of those could also be written as "4l-6"? By the way, I can't give answers, just solutions.

OpenStudy (anonymous):

By that I mean "70 = 4l - 6"

OpenStudy (tatumlee):

I would be upset if you gave me the answer, not what i'm looking for. And I guess i just forgot how to solve an equation I'm very rusty on algebra and missed many steps

OpenStudy (anonymous):

Ok. Do you get it now or do you still need help?

OpenStudy (tatumlee):

so I just do 2* l and 2* 3 wouldn't that be 2l + 2l -6 so 4l-6 and would this be equal to 2l-3 so 70=2l-3?

OpenStudy (anonymous):

4l - 6 isn't equal to 2l - 3, although I can see where you came up with that, dividing everything by 2. However, plug a number like 3 into "l". 4(3) - 6 = 6, while 2(3) - 3 = 3. However, you have stumbled accidentally on part of the answer. The number 2. What is 2 times 2l - 3? Do any of the answers look like that?

OpenStudy (anonymous):

I won't say yes or no to answers, though, so please don't ask for confirmation, kinda puts me in an awkward spot...

OpenStudy (tatumlee):

Once again, Catharsic. I am not looking for an answer from you. I am trying to learn how to do the problem. But thank you,I will find help somewhere else.

OpenStudy (anonymous):

Ok, good luck on that then.